OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, -2) with a slope of 3/5 ? 5x - 3y = 19 5x - 3y = -19 3x - 5y = 19 3x - 5y = -19

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Yah there?

OpenStudy (anonymous):

can you answer this question^

OpenStudy (anonymous):

Ok so we know the form y=mx+b where y and x are you interval values and x is you slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

m is your slope i mean

OpenStudy (anonymous):

so lets plug them in

OpenStudy (anonymous):

-2=3/5(3)+b right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=-2 x=3 m= 3/5

OpenStudy (anonymous):

Then we solve for b

OpenStudy (anonymous):

we find b = 3.8

OpenStudy (anonymous):

plug b into our equation giving us y=3/5x+3.8

OpenStudy (anonymous):

multiply both sides by 5 giving us 5y=3x+19

OpenStudy (anonymous):

subtract 3x over giving 5y-3x=19

OpenStudy (anonymous):

that is the same as 3x-5y=-19

OpenStudy (anonymous):

i believe that is one of your answers

OpenStudy (anonymous):

good?